six carpenters can do a job in nine days. working at the same rate,
1) How many days will it take two carpenters to do the job?
2) How many carpenters will it take to so the job in three days?

Answer by ankor@dixie-net.com(22740)

six carpenters can do a job in nine days. working at the same rate,
:
Find the number of man-days required to complete the job
6*9 = 54 man-days
:
1) How many days will it take two carpenters to do the job?
2d = 54
d = 54/2
d = 27 days
:
2) How many carpenters will it take to so the job in three days?
3m = 54
m = 54/3
m = 18 men